OpenStudy (anonymous):

(x^2-y^2/y^2)/(x+y/y)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{\frac{x^2-y^2}{y^2}}{\frac{x+y}{y}}\] \[x^2-y^2=(x+y)(x-y)\] \[\frac{\cancel{(x+y)}(x-y)}{y^\cancel2}*\frac{\cancel{y}}{\cancel{x+y}}\] \[\frac{x-y}{y}\]